### v2

# InstructTA: Instruction-Tuned Targeted Attack for Large Vision-Language Models

Abstract
Large vision-language models (LVLMs) have demonstrated their incredible capability in image understanding and response generation. However, this rich visual interaction also makes LVLMs vulnerable to adversarial examples. In this paper, we formulate a novel and practical gray-box attack scenario that the adversary can only access the visual encoder of the victim LVLM, without the knowledge of its prompts (which are often proprietary for service providers and not publicly available) and its underlying large language model (LLM). This practical setting poses challenges to the cross-prompt and cross-model transferability of targeted adversarial attack, which aims to confuse the LVLM to output a response that is semantically similar to the attacker’s chosen target text. To this end, we propose an instruction-tuned targeted attack (dubbed InstructTA) to deliver the targeted adversarial attack on LVLMs with high transferability. Initially, we utilize a public text-to-image generative model to “reverse” the target response into a target image, and employ GPT-4 to infer a reasonable instruction $\boldsymbol{p}^{\prime}$ from the target response. We then form a local surrogate model (sharing the same visual encoder with the victim LVLM) to extract instruction-aware features of an adversarial image example and the target image, and minimize the distance between these two features to optimize the adversarial example. To further improve the transferability, we augment the instruction $\boldsymbol{p}^{\prime}$ with instructions paraphrased from an LLM. Extensive experiments demonstrate the superiority of our proposed method in targeted attack performance and transferability.
1 Introduction
Large vision-language models (LVLMs) have achieved great success in text-to-image generation [35], image captioning [7] and visual question-answering [25], due to the colossal increase of training data and model parameters. Benefiting from the strong comprehension of large language models (LLMs), recent LVLMs [25, 12, 48, 27] on top of LLMs have demonstrated a superior capability in solving complex tasks than that of small vision-language models. It is noted that GPT-4 [32] supports visual inputs and perform human-like conversations in various complicated scenarios.
However, recent studies [46, 15, 36, 3] exposed that LLMs with visual structures could be easily fooled by adversarial examples (AEs) crafted via adding imperceptible perturbations to benign images. In particular, targeted attack with malicious goal raises more serious safety concerns to LVLMs and downstream applications. For example, targeted adversarial examples could jailbreak LVLMs to explain the bomb-making process or trick medical diagnostic models into generating reports of specific diseases for insurance. It is thus demanding to comprehensively study adversarial attacks on LVLMs to address these security threats before security breaches happen.
This paper focuses on the targeted attack on LVLMs, where the adversary aims to generate an adversarial image that misleads the LVLM to respond with a specific target text, despite varying instructions the LVLM may use. We target a practical gray-box setting, where the adversary can only access the visual encoder of the LVLM without any knowledge of the LLM and instructions leveraged by the LVLM. [^1] To clarify, in this paper, we use the term “instruction” and “prompt” interchangeably. While de facto LVLMs often leverage standard visual encoders that are publicly known, the employed instructions are often proprietary for service providers and not publicly available. This practical assumption brings some key challenges: (1) The attacker cannot easily gain the instruction-aware features from the accessed visual encoder to launch exact targeted attacks due to a lack of instructions. (2) Cross-model and cross-instruction transferability places new demands on the targeted attack, since the target LVLM may have distinct LLM backends and use varying (unknown) instructions.
Figure 1: The framework of our instruction-tuned targeted attack (InstructTA). Given a target text $\boldsymbol{y}_{t}$, we first transform it into the target image $\boldsymbol{x}_{t}$ with a text-to-image model $h_{\xi}$. Simultaneously, GPT-4 infers a reasonable instruction $\boldsymbol{p}^{\prime}$. When feeding the augmented instruction $\boldsymbol{p}_{i}^{\prime}$ and $\boldsymbol{p}_{j}^{\prime}$ rephrased by an LLM $\boldsymbol{G}_{3.5}$, the surrogate model $M$ extracts instruction-aware features of the AE $\boldsymbol{x}^{\prime}$ and $\boldsymbol{x}_{t}$, respectively. Finally, we minimize the $L_{2}$ distance between these two features to optimize the AE $\boldsymbol{x}^{\prime}$. Notably, the instruction $\boldsymbol{p}_{i/j}^{\prime}$ is updated on-the-fly during the optimization of the adversarial sample.
To overcome these challenges, this paper proposes an instruction-tuned targeted attack (InstructTA) for LVLMs with high attack transferability. As depicted in Fig. 1, we first employ GPT-4 [32] to infer an instruction $\boldsymbol{p}^{\prime}$ in line with the attacker-decided response, and also generate a target image for the target response using a text-to-image generative model. Subsequently, we take both images and instructions as inputs to a local surrogate model to extract informative features tailored to the target response. InstructTA then minimizes the feature distance between the adversarial image example and the target image. Moreover, InstructTA rephrases $\boldsymbol{p}^{\prime}$ into a set of instructions with close semantics, and leverages them to effectively improve the transferability of generated adversarial samples. Our main contributions are outlined as follows:
  • This paper explores the attack vector of LVLMs under a practical yet unexplored gray-box setting, i.e., the adversary only has the rights to access the visual encoder in LVLMs, excluding the LLM and instructions.
  • We propose an effective targeted attack, InstructTA. It cleverly leverages text-to-image models and LLMs to infer reasonable target images and instructions, respectively, and augments the transferability of adversarial samples with instruction-tuning paradigms.
  • Extensive experiments demonstrate that InstructTA can achieve better targeted attack performance and transferability than intuitive baselines. We also discuss mitigations and ethics considerations of InstructTA.
2 Related Work
LVLMs. Benefiting from the powerful capabilities of LLMs in answering diverse and complex questions [31], LVLMs [46] have demonstrated remarkable performance on a variety of multi-modal tasks, e.g., image captioning, visual question-answering, natural language visual reasoning and visual dialog. To bridge the image-text modality gap, BLIP-2 [25] proposes the Q-Former to extract visual features into the frozen OPT [45] or FlanT5 [11]. MiniGPT-4 [48] employs the identical pretrained ViT [16] and Q-Former as BLIP-2, but uses the more powerful Vicuna [10] to serve as the base LLM for text generation. InstructBLIP [12] introduces an instruction-aware Q-Former to extract the most task-relevant visual representations. LLaVA [27] bridges a pre-trained visual encoder with LLaMA, and fine-tunes them with question-image-answer paired data generated by GPT-4.
Figure 2: The architecture of LVLM.
Adversarial Attacks. Although adversarial attacks are now being developed in a variety of modalities and tasks, the earliest and most extensive research occurred in the field of image classification [4, 37]. In this case, the goal of adversarial attack is often to fool a well-trained visual classifier by crafted adversarial images. Based on their purpose, we divide adversarial attacks into targeted attacks and non-targeted attacks [2]. Targeted attacks aim to confuse the model to recognize the adversarial example as a target label, whereas non-targeted attacks only ensure that the model make wrong predictions, not caring about the output error category. Moreover, according to the information of the attacked model exposed to the adversary, adversarial attacks can be grouped into white-box attack and black-box attack. The white-box attacks can obtain the whole parameters and structure of the target model. Thus, they generally optimize adversarial perturbations with gradients of loss w.r.t inputs, e.g., FGSM [19], I-FGSM [22], C&W[5] and PGD [30]. In contrast, only the input and output of the attacked model can be captured by black-box attack, which is more challenging and practical than white-box one. One solution is to estimate the gradient or search for successful perturbations by directly querying the target model multiple times, e.g., ZOO [8] and NES attack [21]. Another way is to implement transfer attacks by constructing adversarial samples from substitute models, e.g., SBA [33] and ensemble-based approaches [28, 14, 18, 20].
Besides image classification, researchers have extended adversarial attacks to vision-language models or tasks [6, 40, 23, 38, 26, 43, 41, 47, 29, 39, 46]. Recently, investigating adversarial robustness on vision-language pre-training (VLP) models [43, 29, 39] has aroused much attention. Zhang et al. [43] proposed Co-Attack for VLP models, which collaboratively combines the image and the text perturbations. SGA [29] improved the adversarial transferability across different VLP models by breaking the set-level cross-modal alignment.
AEs on LVLMs. Currently, there are some pioneering and contemporaneous works of targeted attacks on LVLMs. On one hand, some works aim to mislead the LVLMs to output specific words. For example, Bailey et al. [3] forces the LVLM to respond with the target string in a white-box setting. Dong et al. [15] present text description attack to mislead the black-box Bard [1] to output targeted responses, which maximize the log-likelihood of the substitute LVLMs to predict the target sentence. On the other hand, some works have been devoted to fooling LVLMs into outputting text with the same semantics as the target. For example, Zhao et al. [46] propose three feature matching methods (i.e., MF-it, MF-ii and MF-tt) to launch targeted attack with given target text against LVLMs in black-box settings. Shayegani et al. [36] also align the adversarial image and target images in the embedding space for LVLM jailbreak. Different from [46], we focus on the gray-box setting instead of the black-box setting, and further improve the attack transferability under the guidance of inferred and rephrased instructions. While [36] also accesses visual encoders in LVLMs for jailbreak, our work highlights the capability to implement generic targeted attacks even without knowing the target LVLM’s instructions.
3 Method
3.1 Problem Formulation
LVLM. Given an image $\boldsymbol{x}$ and a text instruction $\boldsymbol{p}$ as inputs, the LVLM $F$ outputs an answer $\boldsymbol{y}$. Formally,
$\boldsymbol{y}=F_{\theta}(\boldsymbol{x},\boldsymbol{p}),$ (1)
where $F_{\theta}$ denotes the LVLM parameterized by $\theta$. Typically, the LVLM [25, 48, 12, 27] projects the extracted visual features into the word embedding space and then feeds them to its underlying LLM. Hence, LVLM consists of a pre-trained visual encoder, a projector and a well-trained LLM, as shown in 2.
Adversarial Objective. In image-grounded text generation, the goal of targeted attack is to construct adversarial examples, which could cause the output of the victim model to match a predefined targeted response. Under the $L_{\infty}$ norm constraint, the objective of a targeted attack can be formulated as follows:
$\Delta(\boldsymbol{x},\boldsymbol{y}_{t}):$ $=\max_{\boldsymbol{x}^{\prime}}S(F(\boldsymbol{x}^{\prime},\boldsymbol{p}),\boldsymbol{y}_{t}),$ (2)
s.t. $\|\boldsymbol{x}^{\prime}-\boldsymbol{x}\|_{\infty}\leq\epsilon,$
where $\boldsymbol{x}$ is a given benign image sample, $\boldsymbol{x}^{\prime}$ represents the adversarial example of $\boldsymbol{x}$, and $\boldsymbol{y}_{t}$ denotes the predefined target text. $\epsilon$ is the budget of controlling adversarial perturbations, and $\boldsymbol{p}$ denotes the instructions used by the target LVLM ($\boldsymbol{p}$ is unknown to attackers; see § 3.2). $S(\cdot,\cdot)$ is a metric depicting the semantic similarity between two input sentences. This way, Eqn. 2 ensures that the response of the adversarial example shall retain the same meaning as the target text $y_{t}$.
Algorithm 1 InstructTA
0:  a benign image $\boldsymbol{x}$, a target text $\boldsymbol{y}_{t}$, a text-to-image generative model $h_{\xi}$, a surrogate model $M$, a VLP model with the visual encoder $f$ and text encoder $g$, ChatGPT $G_{3.5}$, GPT-4 $G_{4}$, custom prompts $\boldsymbol{p}_{g}$ and $\boldsymbol{p}_{r}$, step size $\eta$, perturbation budget $\epsilon$, attack iterations $T$
1:  Initialize: $\boldsymbol{x}^{\prime}=\boldsymbol{x}$, $\boldsymbol{x}_{t}=h_{\xi}(\boldsymbol{y}_{t})$, $\boldsymbol{p^{\prime}}=G_{4}(\boldsymbol{p}_{g}\boldsymbol{y}_{t})$  // Eqn. (3)
2:  for $iter=1...T$ do
3:     $\boldsymbol{p}_{i}^{\prime}\leftarrow G_{3.5}(\boldsymbol{p}_{r}\boldsymbol{p}^{\prime})$
4:     $\boldsymbol{p}_{j}^{\prime}\leftarrow G_{3.5}(\boldsymbol{p}_{r}\boldsymbol{p}^{\prime})$
5:     $\boldsymbol{x}^{\prime}={\mathcal{S}}_{\epsilon}(\boldsymbol{x}^{\prime}-\eta\cdot\operatorname{sign}(\Delta_{\boldsymbol{x}^{\prime}}\mathcal{L}))$  // PGD
6:  end for
6:  adversarial example $\boldsymbol{x}^{\prime}$
Figure 3: An example of rephrasing an instruction. Given the target response $\boldsymbol{y}_{t}$, GPT-4 infers an instruction, i.e., “Can you describe the scene depicted in the photograph?”. The real instruction assigned to this target text is “Explain the visual content of the image in great detail.”
3.2 Targeted Attack with Instruction Tuning
Assumption. In this paper, we consider a practical gray-box setting, where the adversary only knows the visual encoder used in the attacked model, excluding LLMs and instructions of the victim LVLM. That is, the adversary is agnostic to $\boldsymbol{p}$. Instead, we assume that attackers know the usage domain of the target LVLM, e.g., CT image diagnosis; considering $F$ is often hosted as a task-specific service, attackers can easily obtain this information. Accordingly, attackers can locally prepare instructions $\boldsymbol{p}^{\prime}$ which are commonly used in the target domain (see below for details). We also assume that before attack, attackers can form a local surrogate model with the same visual encoder as that of the victim model.
Inferring Instructions. We first locally infer the possible instructions $\boldsymbol{p}^{\prime}$ for the target text $\boldsymbol{y}_{t}$ by using GPT-4 [32]. Formally,
$\boldsymbol{p^{\prime}}=G_{4}(\boldsymbol{p}_{g}\boldsymbol{y}_{t}),$ (3)
where $G_{4}$ denotes the language-only GPT-4, and $\boldsymbol{p}^{\prime}$ represents an instruction generated by $G_{4}$. The above action denotes that we feed an instruction $\boldsymbol{p}_{g}$ concatenated with the target response text $\boldsymbol{y}_{t}$ to the GPT-4. In the experiment, we set $\boldsymbol{p}_{g}$ to a straightforward prompt “What is the question or prompt for this answer: ”.
Forming Surrogate Models. Once the instructions have been inferred, the surrogate model uses a projector to transform the features from the visual encoder into instruction-aware features. Thus, the surrogate model comprises two modules: the visual encoder from the victim model and the pre-trained projector (e.g., Q-Former [25]). The projector maps global visual features into informative features tailored to given instructions.
Generating Adversarial Examples. We then extract the instruction-aware visual features most conducive to the task by inputting both images and the generated instructions to the surrogate model. For the targeted attack, a simple solution is to align the adversarial example to the target text in the feature space of the substitute model. Given that the surrogate model only extracts visual features, it fails to build target embedding for the target text. An alternative approach is to transform the target text into the target image through text-to-image generative models (e.g., Stable Diffusion [35]). Hence, we reduce the embedding distance between the adversarial example and the target image $\boldsymbol{x}_{t}$ for the targeted attack as:
$\min_{\boldsymbol{x}^{\prime}}\|M(\boldsymbol{x}^{\prime},\boldsymbol{p}^{\prime})-M(h_{\xi}(\boldsymbol{y}_{t}),\boldsymbol{p}^{\prime})\|_{2}^{2},{\quad\text{s.t.}~{}}\|\boldsymbol{x}^{\prime}-\boldsymbol{x}\|_{\infty}\leq\epsilon$ (4)
where $M(\cdot,\cdot)$ denotes the substitute model, $h_{\xi}$ is a public generative model ($\boldsymbol{x}_{t}=h_{\xi}(\boldsymbol{y}_{t})$), and $\boldsymbol{x}^{\prime}$ indicates the adversarial sample of $\boldsymbol{x}$.
Augmenting Transferability. Our tentative exploration shows that AEs generated by Eqn. 4, although manifesting reasonable attackability, are often less transferable to the wide range of possible instructions in the wild. Thus, to improve the transferability and robustness of the generated adversarial examples, we use ChatGPT (i.e., GPT-3.5) to rephrase the inferred $\boldsymbol{p}^{\prime}$ as augmentation. The paraphrase prompt template fed into the ChatGPT is “Paraphrase this sentence: ”. This way, Eqn. 4 can be rewritten as
$\min_{\boldsymbol{x}^{\prime}}\mathcal{J}_{ins}=\mathbb{E}_{\boldsymbol{p}_{i}^{\prime},\boldsymbol{p}_{j}^{\prime}}\|M(\boldsymbol{x}^{\prime}$ $,\boldsymbol{p}_{i}^{\prime})-M(h_{\xi}(\boldsymbol{y}_{t}),\boldsymbol{p}_{j}^{\prime})\|_{2}^{2},$ (5)
[…]
where $\boldsymbol{p}_{i}^{\prime}$ and $\boldsymbol{p}_{j}^{\prime}$ are augmented instructions from $\boldsymbol{p}^{\prime}$. Both of them are generated by ChatGPT, i.e., $\boldsymbol{p}_{i}^{\prime}=G_{3.5}(\boldsymbol{p}_{r}\boldsymbol{p}^{\prime})$, $\boldsymbol{p}_{j}^{\prime}=G_{3.5}(\boldsymbol{p}_{r}\boldsymbol{p}^{\prime})$, where $G_{3.5}$ represents ChatGPT and $\boldsymbol{p}_{r}$ denotes the aforementioned rephrasing template. As illustrated in Fig. 3, ChatGPT is generally effective to rephrase a seed instruction $\boldsymbol{p}^{\prime}$ into a set of semantically-correlated instructions.
Table 1: Targeted attacks against victim LVLMs. A total of 1,000 clean images denoted as $\boldsymbol{x}$ are sampled from the ImageNet-1K [13] validation set. For each of these benign images, a targeted text $\boldsymbol{y}_{t}$ and a prompt $\boldsymbol{p}$ are randomly assigned from LLaVA-Instruct-150K [27]. Our analysis involves the computation of the CLIP score (normal-$\uparrow$) [34, 46] between the target texts and the generated responses of AEs on attacked LVLMs, which was performed using several CLIP text encoders and their ensemble. We also report the No. of successful attacks within 1,000 adversarial samples, determined by GPT-4. CLIP and EVA-CLIP in parentheses indicate that they are used as surrogate models for transfer attack.
VLVM Attack method Text encoder (pretrained) for evaluation The No. of successes RN50 RN101 ViT-B/16 ViT-B/32 ViT-L/14 Ensemble BLIP-2 Clean image 0.296 0.427 0.342 0.354 0.171 0.318 4 MF-it (CLIP) [46] 0.292 0.424 0.339 0.353 0.168 0.315 2 MF-ii (CLIP) [46] 0.293 0.426 0.340 0.353 0.168 0.316 2 MF-it (EVA-CLIP) 0.574 0.667 0.616 0.630 0.481 0.594 519 MF-ii (EVA-CLIP) 0.573 0.661 0.616 0.627 0.484 0.592 514 InstructTA-MF-it (ours) 0.604 0.691 0.644 0.656 0.515 0.622 652 InstructTA-MF-ii (ours) 0.599 0.682 0.639 0.649 0.512 0.616 619 InstructBLIP Clean image 0.392 0.557 0.420 0.421 0.312 0.420 0 MF-it (CLIP) [46] 0.394 0.558 0.423 0.422 0.313 0.422 0 MF-ii (CLIP) [46] 0.395 0.557 0.424 0.423 0.314 0.422 0 MF-it (EVA-CLIP) 0.773 0.822 0.791 0.796 0.740 0.785 292 MF-ii (EVA-CLIP) 0.761 0.813 0.780 0.783 0.727 0.773 259 InstructTA-MF-it (ours) 0.810 0.849 0.826 0.830 0.784 0.820 383 InstructTA-MF-ii (ours) 0.789 0.833 0.805 0.809 0.758 0.799 328 MiniGPT-4 Clean image 0.345 0.495 0.367 0.374 0.253 0.367 0 MF-it [46] 0.341 0.493 0.364 0.369 0.251 0.364 0 MF-ii [46] 0.340 0.492 0.363 0.369 0.249 0.362 2 MF-it (EVA-CLIP) 0.650 0.726 0.664 0.676 0.593 0.662 259 MF-ii (EVA-CLIP) 0.649 0.723 0.665 0.670 0.592 0.660 247 InstructTA-MF-it (ours) 0.692 0.756 0.701 0.712 0.638 0.700 363 InstructTA-MF-ii (ours) 0.680 0.748 0.692 0.701 0.627 0.690 325
3.3 Dual Targeted Attack & Optimization
With the aforementioned steps, we form following dual targeted attack framework:
$\arg\min_{\boldsymbol{x}^{\prime}}\mathcal{L}=\mathcal{J}_{ins}-\mathcal{J}_{mf},{\quad\text{s.t.}~{}}\|\boldsymbol{x}^{\prime}&-\boldsymbol{x}\|_{\infty}\leq\epsilon\end{aligned},$ (6)
where $\mathcal{J}_{mf}$ is the objective of MF-it or MF-ii, following [46]. MF-it and MF-ii achieve transfer targeted attack via matching the embeddings of the adversarial example and the target response. For MF-it, its objective is defined as
[…]
where $f(\cdot)$ and $g(\cdot)$ are the paired visual encoder and text encoder of an align-based VLP model (e.g., CLIP [34], ALBEF [24], Open-CLIP [9], and EVA-CLIP [17]), respectively. For MF-ii, we have
[…]
Notably, Eqn. 6 with MF-it is marked InstructTA-MF-it, and Eqn. 6 with MF-ii is marked InstructTA-MF-ii.
When optimizing adversarial samples following Eqn. 6, this paper adopts PGD to update $\boldsymbol{x}^{\prime}$ with $T$ iterations as below,
$\boldsymbol{x}_{T}^{\prime}={\mathcal{S}}_{\epsilon}(\boldsymbol{x}_{T-1}^{\prime}-\eta\cdot\operatorname{sign}(\Delta_{\boldsymbol{x}_{T-1}^{\prime}}\mathcal{L})),\quad\boldsymbol{x}^{\prime}_{0}=\boldsymbol{x},$ (9)
where $\eta$ denotes step size, and $\mathcal{S}_{\epsilon}$ clips $\boldsymbol{x}^{\prime}$ into the $\epsilon$-ball [30] of $\boldsymbol{x}$.
In Algorithm 1, we outline the whole pipeline of InstructTA. Initially, once decided an adversarial target text $\boldsymbol{y}_{t}$, we use the text-to-image model $h_{\xi}$ and GPT-4 to “reverse” a target image $\boldsymbol{x}_{t}$ and produce the reference instruction $\boldsymbol{p}^{\prime}$, respectively. In each iteration, we augment the inferred instruction $\boldsymbol{p}^{\prime}$ with ChatGPT firstly. Then, we update the adversarial example with PGD over Eqn. 6. Until the end of $T$ step iterations, the final adversarial sample is obtained and used for attacking remote LVLMs.
4 Experiments
This section presents the experimental results of the proposed InstructTA on multiple victim LVLMs. We introduce the evaluation setup in § 4.1, and then present the results in § 4.2 followed by ablation study in § 4.3 and discussion in § 4.4.
4.1 Evaluation Setup
Datasets. The ImageNet-1K [13] comprises 3 subsets, including training, validation and test sets. We randomly select 1,000 images from the validation set to serve as the benign samples ($\boldsymbol{x}$) for the attack. Correspondingly, we sample 1,000 instruction-answer pairs from LLaVA-Instruct-150K [27] as tested instructions and target texts, respectively, since LLaVA-Instruct-150K is a set of multimodal instruction-following data generated by GPT-4. In addition, we use Stable Diffusion ($h_{\xi}$) to generate target images from the target texts, following [46].
Protocols. Following [46], we adopt CLIP-score ($\uparrow$) to evaluate the targeted attack performance, which calculates the semantic similarity between the generated response of victim LVLM and the target text using CLIP text encoders. Additionally, we record the number of successful attacks by using GPT-4 to determine whether two texts have the same meaning. The prompt in this operation is as follows: Determine whether these two texts describe the same objects or things, you only need to answer yes or no: 1. $\boldsymbol{y}^{\prime}$ 2. $\boldsymbol{y}_{t}$ where $\boldsymbol{y}^{\prime}$ denotes the output response of the adversarial example on victim model, and $\boldsymbol{y}_{t}$ is the target response.
Baselines. This paper provides evaluations on three victim LVLMs, including BLIP-2 [25], InstructBLIP [12] and MiniGPT-4 [48]. In detail, we select flan-t5-xxl [^2] https://huggingface.co/google/flan-t5-xxl, vicuna-7b-v1.1333https://huggingface.co/lmsys/vicuna-7b-v1.1 and vicuna-v0-13b [^4] https://huggingface.co/Vision-CAIR/vicuna as LLMs for BLIP-2, InstructBLIP and MiniGPT-4, respectively. Their visual encoders are all ViT-G/14 [42] borrowed from EVA-CLIP [17]. To evaluate the targeted attack performance, InstructTA is compared with existing baselines, i.e., MF-it and MF-ii. They can choose CLIP (ViT-B/32) [34] or the EVA-CLIP (ViT-G/14) as surrogate models for targeted attacks.
Implementation Details. All images are resized to $224\times 224$ before feeding in substitute models and LVLMs. For InstructTA, we choose the visual encoder and Q-Former from InstructBLIP as the surrogate model. For PGD, $\eta$ and $T$ are set to 1 and 100, respectively. The maximum perturbation magnitude $\epsilon$ is set to $8$. We implement all the evaluations using Pytorch 2.1.0 and run them on a single NVIDIA RTX A6000 GPU (48G).
(a) The target text and image
(b) Clean image
(c) MF-it (CLIP)
(d) MF-ii (CLIP)
(e) MF-it (EVA-CLIP)
(f) MF-ii (EVA-CLIP)
(g) InstructTA-MF-it
(h) InstructTA-MF-ii
Figure 4: Visualization examples of various targeted attack methods on InstructBLIP. “Explain the visual content of the image in great detail.” is a real instruction.
4.2 Results
Targeted Attack Performance. As illustrated in Table 1, we transfer adversarial samples generated by different targeted attack methods to victim LVLMs and mislead them to produce target responses. For each adversarial example, when attacking a LVLM, the target text and the assigned instruction may be different from other adversarial examples, since we randomly selected 1000 pairs of target texts and instructions from LLaVA-Instruct-150K [27]. Subsequently, we evaluate the similarity between the generated response $\boldsymbol{y}^{\prime}$ of the victim LVLM and the targeted text $\boldsymbol{y}_{t}$ by utilizing diverse forms of CLIP text encoders, where $\boldsymbol{y}^{\prime}=F(\boldsymbol{x}^{\prime},\boldsymbol{p})$. In addition, we count the number of successful targeted adversarial samples, as judged by GPT-4.
Table 1 reports the evaluation results of different targeted attack methods on BLIP-2, InstructBLIP and MiniGPT-4. We compare InstructTA with MF-it and MF-ii, which are the two recent attacks on LVLMs. Overall, we observe that InstructTA consistently outperforms MF-it and MF-ii in terms of CLIP score and the number of successful attacks. In fact, it can be seen that MF-it and MF-ii with CLIP have almost the same results as the clean samples, indicating no targeted attack effect. We believe this is reasonable, because ViT-B/32 in CLIP shares a different visual encoder than that of the target LVLM. In contrast, MF-it (EVA-CLIP), MF-ii (EVA-CLIP) and our methods all notably outperform the targeted attack results on clean samples because they share the same visual encoder as the victim model. Overall, we interpret these results justify the necessity of using the same visual encoder as the victim model.
More importantly, both InstructTA-MF-it and InstructTA-MF-ii are distinctly better than MF-it and MF-ii in terms of CLIP-score and the number of successful attacks. These phenomena demonstrate that our method can effectively improve the transferability of the targeted attack on LVLMs. One reason is that we enhance the robustness of AEs to varying instructions. Another reason is that our dual attacks (InstructTA-MF-it/InstructTA-MF-ii) extracts informative features containing global and instruction-aware knowledges, benefiting to precise targeted attacks.
Table 2: Ablation study of the proposed method on BLIP-2. CLIP score [46] is computed by ensemble CLIP text encoders. Attack success rate is the ratio between the No. of attack successes and the total No. of samples.
Attack method CLIP score ($\uparrow$) Attack success rate ($\uparrow$) InstructTA-MF-it 0.622 0.652 InstructTA-MF-ii 0.616 0.619 InstructTA 0.606 0.624 InstructTA-woG 0.571 0.448
Visualization Examples. Fig. 3 has displayed a sample inferred instruction and its sibling instructions rephrased by ChatGPT. It can be seen that the instructions inferred by GPT-4 are similar to the real instruction. The rephrased instructions are even more similar to the original sentence.
Moreover, we also provide some visual results of different attack methods on attacking InstructBLIP, as shown in Fig. 4. Overall, it can be observed that the generated responses of our methods are closer to the target texts.
Table 3: Targeted attack performance for multiple perturbation budget $\epsilon$.
Budget $\epsilon$ Text encoder (pretrained) for evaluation Attack success rate RN50 RN101 ViT-B/16 ViT-B/32 ViT-L/14 Ensemble 2 0.437 0.547 0.480 0.492 0.330 0.457 0.154 4 0.560 0.651 0.602 0.612 0.470 0.579 0.450 8 0.604 0.691 0.644 0.656 0.515 0.622 0.652 16 0.617 0.702 0.657 0.669 0.527 0.634 0.746 64 0.617 0.704 0.658 0.671 0.529 0.636 0.740
Figure 5: To explore the impact of varying $\epsilon$ values within the InstructTA-MF-it, we conducted experiments aiming to achieve different levels of perturbed images on BLIP-2, i.e., referred to as the AE $\boldsymbol{x}^{\prime}$. Our findings indicate a degradation in the visual quality of $\boldsymbol{x}^{\prime}$, as quantified by the LPIPS [44] distance between the original image $\boldsymbol{x}$ and the adversarial image $\boldsymbol{x}^{\prime}$. Simultaneously, the effectiveness of targeted response generation reaches a saturation point. Consequently, it is crucial to establish an appropriate perturbation budget, such as $\epsilon=8$, to effectively balance the image quality and the targeted attack performance.
4.3 Ablation Study
We also launched ablation study to explore the contribution of major components in the proposed method. At this step, we conducted experiments on BLIP-2 with two key component ablated sequentially: 1) removing MF-it and MF-ii (denoted as InstructTA), and 2) launching optimization without CLIP embedding matching nor ChatGPT for paraphrase (in accordance with Eqn. 4; denoted as InstructTA-woG).
The results are reported in Table 2. It is observed that the performances degrade when one or more parts are removed. In particular, the InstructTA-woG setting exhibits a success rate downgrading of over 31% in comparison to the InstructTA-MF-it setting. Overall, we interpret the ablation study results as reasonable, illustrating that all the constituent components have been thoughtfully devised and exhibit congruity, culminating in an efficacious solution for the targeted attack.
4.4 Discussion
Effect of $\epsilon$. To explore the effect of the perturbation magnitude $\epsilon$ on targeted attack performance, we study the attack results of InstructTA-MF-it toward BLIP-2 under various $\epsilon$. The results are reported in Table 3. It is evident that with $\epsilon$ increasing, both CLIP scores and attack success rates rise, illustrating stronger attack performance. Nevertheless, it is seen that the attack performance becomes gradually saturated when $\epsilon$ is greater than $16$. Moreover, as shown in Fig. 5, an overly large $\epsilon$ (e.g., 16 and 64) has made the perturbation perceptible to naked eyes. Thus, we suggest to explore a proper $\epsilon$ for the targeted attack, e.g., $\epsilon=8$, such that the attack performance can be effectively improved while the perturbation is imperceptible.
Table 4: Average CLIP score ($\uparrow$) between different versions of instructions. Each type of instruction have 1,000 samples.
Instruction Rephrased instruction $\boldsymbol{p}_{i/j}^{\prime}$ Real instruction $\boldsymbol{p}$ Inferred instruction $\boldsymbol{p}^{\prime}$ 0.942 0.883 Real instruction $\boldsymbol{p}$ 0.882 1.000
Assessment of Generated Instructions. To quantitatively evaluate the usability of the generated instructions, we collect a paraphrased instruction for each inferred instruction used in main experiment and provide average CLIP score values between them, as reported in Table 4. It is evident that there exists a high similarity score between the rephrased and the inferred instructions; moreover, the rephrased and the inferred instructions both show high similarity with the real instructions. We interpret the results as encouraging, demonstrating the high quality of employed instructions.
5 Conclusion, Ethics, and Mitigation
Our work highlights the vulnerability of LVLMs to adversarial attacks under a practical, gray-box scenario. InstructTA delivers targeted attacks on LVLMs with high transferability, demonstrating the importance of instruction-based techniques and the fusion of visual and language information in boosting adversarial attacks. The aim of our work is to enhance the resilience of LVLMs. Ethically, since we discovered adversarial examples, we minimized harm by (1) avoiding any damage to real users and (2) disclosing the detailed attacking process and methods used to find them in this paper. Overall, we believe that the security of the LVLM ecosystem is best advanced by responsible researchers investigating these problems. Looking ahead, we envision the feasibility of launching various mitigation methods to improve the robustness of LVLMs against adversarial attacks. For instance, we can leverage the adversarial training framework [30] to train LVLMs with adversarial examples. We can also explore the possibility of incorporating adversarial training into the pre-training stage of LVLMs or their visual encoders. Moreover, we can also explore detecting adversarial examples by leveraging the discrepancy between the instruction-aware features of the adversarial examples and the original images.
Supplementary Material
6 Attack Cases
Fig. 6 shows an example to attack BLIP-2 for different methods. It can be seen from the figure that CLIP-based MF-it and MF-ii shows no targeted attack effect since they have the similar results with the clean sample. Interestingly, half of the response (i.e., “a man and a woman”) of the adversarial sample generated by MF-it (EVA-CLIP) matches the original image and half (i.e., “elephants in a zoo”) matches the target image. MF-ii (EVA-CLIP), InstructTA-MF-it and InstructTA-MF-ii demonstrate the targeted attack effect. Importantly, the description of our InstructTA-MF-ii best matches the target image.
7 Results with Complex Reasoning Questions
Inferring from images can demonstrate the model’s deep understanding of the image semantics. The box below shows some reasoning instructions from complex_reasoning_77k in LLaVA-Instruct-150K [27]. 1. How can pedestrians ensure their safety when navigating this busy city scene at night? 2. What might the cat be curious about on the floor? 3. What suggestions can be made for individuals wishing to serve these dough balls in an aesthetically pleasing manner? 4. What might be the man’s intention in the kitchen? 5. How does the position of the players on the field affect their performance? 6. What unique feature can be observed about the desserts in the scene? 7. What type of event might be taking place in this scene? 8. What could be a possible reason for the dog and cat’s interaction? 9. In terms of architectural features, why might this clock tower be studied or admired? 10. What tactic might the man be employing during the tennis match? 11. What could be the significance or purpose of a single pink rose placed in a glass with water? 12. What can you infer about the laptop owner’s personal style or interests? 13. How is the truck being used during this event, and what challenges might it face in this context? 14. Why might the vandalized stop sign on the school bus cause concern? …
To further evaluate the effectiveness of the proposed methods, we provide transfer attack results for complex reasoning instructions, as portrayed in Table 6. All methods, even clean samples, have obvious targeted attack performance. This may be because some of the selected reasoning instructions can be answered without the need for image information, e.g., “Why might the vandalized stop sign on the school bus cause concern?”. Importantly, our methods have better transfer attack performances in terms of both CLIP score and the No. of successes. It implies that InstructTA still performs well even when faced with complex inference prompts.
8 Victim Models
Table 5 shows the architectures of the target LVLMs.
Table 5: The architectures of the victim LVLMs.
LVLM Visual Encoder Projector LLM BLIP-2 ViT-G/14 from EVA-CLIP Q-Former & Single linear layer Flan-T5-XXL InstructBLIP ViT-G/14 from EVA-CLIP Q-Former & Single linear layer Vicuna-7b-v1.1 MiniGPT-4 ViT-G/14 from EVA-CLIP Q-Former & Single linear layer Vicuna-v0-13b
Figure 6: An instance of different targeted attack methods on BLIP-2. The real instruction is “Can you describe the main features of this image for me?”. The target text is “The image depicts a large elephant and a smaller elephant standing in a grassy area next to a stone wall, likely within a zoo enclosure. The two elephants appear to be gathered near a flock of black birds. The birds can be seen scattered around the enclosure, with some standing near the elephants and others in the background. The scene captures the interaction between the elephants and the birds, creating a serene and captivating moment.”
Table 6: Targeted attacks against victim BLIP-2 with reasoning instructions. We randomly select a total of 1,000 clean images from the ImageNet-1K [13] validation set. For each of these benign sample, a target text and a prompt are randomly assigned from complex_reasoning_77k in LLaVA-Instruct-150K [27]. Our metrics involves the computation of the CLIP score (normal-$\uparrow$) [34, 46] between the target texts and the generated responses of AEs on the attacked LVLM, which was performed using several CLIP text encoders and their ensemble. We also report the No. of successful attacks within 1,000 adversarial samples, determined by GPT-4. CLIP and EVA-CLIP in parentheses indicate that they are used as surrogate models for transfer attack.
VLVM Attack method Text encoder (pretrained) for evaluation The No. of successes RN50 RN101 ViT-B/16 ViT-B/32 ViT-L/14 Ensemble BLIP-2 Clean image 0.609 0.662 0.646 0.651 0.506 0.615 381 MF-it (CLIP) [46] 0.606 0.658 0.643 0.648 0.503 0.612 370 MF-ii (CLIP) [46] 0.606 0.658 0.642 0.648 0.502 0.611 377 MF-it (EVA-CLIP) 0.650 0.698 0.684 0.689 0.554 0.655 547 MF-ii (EVA-CLIP) 0.646 0.696 0.679 0.685 0.545 0.650 529 InstructTA-MF-it (ours) 0.650 0.699 0.683 0.689 0.551 0.654 566 InstructTA-MF-ii (ours) 0.652 0.699 0.685 0.692 0.552 0.656 571

### v3

@@
 However, this rich visual interaction also makes LVLMs vulnerable to adversarial examples.
-In this paper, we formulate a novel and practical gray-box attack scenario that the adversary can only access the visual encoder of the victim LVLM, without the knowledge of its prompts (which are often proprietary for service providers and not publicly available) and its underlying large language model (LLM).
+In this paper, we formulate a novel and practical targeted attack scenario that the adversary can only know the vision encoder of the victim LVLM, without the knowledge of its prompts (which are often proprietary for service providers and not publicly available) and its underlying large language model (LLM).
 This practical setting poses challenges to the cross-prompt and cross-model transferability of targeted adversarial attack, which aims to confuse the LVLM to output a response that is semantically similar to the attacker’s chosen target text.
@@
 Initially, we utilize a public text-to-image generative model to “reverse” the target response into a target image, and employ GPT-4 to infer a reasonable instruction $\boldsymbol{p}^{\prime}$ from the target response.
-We then form a local surrogate model (sharing the same visual encoder with the victim LVLM) to extract instruction-aware features of an adversarial image example and the target image, and minimize the distance between these two features to optimize the adversarial example.
-To further improve the transferability, we augment the instruction $\boldsymbol{p}^{\prime}$ with instructions paraphrased from an LLM.
+We then form a local surrogate model (sharing the same vision encoder with the victim LVLM) to extract instruction-aware features of an adversarial image example and the target image, and minimize the distance between these two features to optimize the adversarial example.
+To further improve the transferability with instruction tuning, we augment the instruction $\boldsymbol{p}^{\prime}$ with instructions paraphrased from GPT-4.
 Extensive experiments demonstrate the superiority of our proposed method in targeted attack performance and transferability.
+The code is available at https://github.com/xunguangwang/InstructTA.
 1 Introduction
-Large vision-language models (LVLMs) have achieved great success in text-to-image generation [35], image captioning [7] and visual question-answering [25], due to the colossal increase of training data and model parameters.
-Benefiting from the strong comprehension of large language models (LLMs), recent LVLMs [25, 12, 48, 27] on top of LLMs have demonstrated a superior capability in solving complex tasks than that of small vision-language models.
-It is noted that GPT-4 [32] supports visual inputs and perform human-like conversations in various complicated scenarios.
+Large vision-language models (LVLMs) have achieved great success in text-to-image generation [41], image captioning [9], visual question-answering [28], visual dialog [21], etc., due to the colossal increase of training data and model parameters.
+Benefiting from the strong comprehension of large language models (LLMs), recent LVLMs [28, 14, 57, 32, 4, 46, 2, 3, 21] on top of LLMs have demonstrated a superior capability in solving complex tasks than that of small vision-language models.
+It is noted that GPT-4 [37] supports visual inputs and performs human-like conversations in various complicated scenarios.
 However, recent studies [55, 16, 42, 5] exposed that LLMs with visual structures could be easily fooled by adversarial examples (AEs) crafted via adding imperceptible perturbations to benign images.
-In particular, targeted attack with malicious goal raises more serious safety concerns to LVLMs and downstream applications.
+In particular, targeted attack with malicious goals raises more serious safety concerns for LVLMs and their downstream applications.
 For example, targeted adversarial examples could jailbreak LVLMs to explain the bomb-making process or trick medical diagnostic models into generating reports of specific diseases for insurance.
-It is thus demanding to comprehensively study adversarial attacks on LVLMs to address these security threats before security breaches happen.
+It is thus demanding to comprehensively study adversarial attacks on LVLMs to address these security threats before safety breaches happen.
 This paper focuses on the targeted attack on LVLMs, where the adversary aims to generate an adversarial image that misleads the LVLM to respond with a specific target text, despite varying instructions the LVLM may use.
-We target a practical gray-box setting, where the adversary can only access the visual encoder of the LVLM without any knowledge of the LLM and instructions leveraged by the LVLM.
-[^1] To clarify, in this paper, we use the term “instruction” and “prompt” interchangeably.
-While de facto LVLMs often leverage standard visual encoders that are publicly known, the employed instructions are often proprietary for service providers and not publicly available.
-This practical assumption brings some key challenges: (1) The attacker cannot easily gain the instruction-aware features from the accessed visual encoder to launch exact targeted attacks due to a lack of instructions.
-(2) Cross-model and cross-instruction transferability places new demands on the targeted attack, since the target LVLM may have distinct LLM backends and use varying (unknown) instructions.
+Typically, LVLMs are deployed by service providers for vertical applications in the cloud, and the built-in prompts [^1] To clarify, in this paper, we use the term “instruction” and “prompt” interchangeably.
+and the parameters of LLMs leveraged by the LVLMs are not publicly accessible as they are proprietary to the service providers.
+Conversely, it is reasonable to assume that attackers possess knowledge of the vision encoder in the LVLM, as developers often disclose the structure of the model in technical reports or attackers can acquire this information through social engineering.
+Consequently, adversaries can easily construct surrogate models with the same vision encoders to facilitate attacks.
+Nonetheless, the uncertainty surrounding the LVLM model itself and its internal instructions presents challenges for targeted attacks: (1) The attacker cannot easily gain the instruction-aware features from the vision encoder to launch exact targeted attacks due to a lack of instructions.
+(2) Cross-model and cross-instruction transferability places new demands on the targeted attack, since the victim LVLM may have distinct LLM backends and use varying (unknown) instructions.
 Figure 1: The framework of our instruction-tuned targeted attack (InstructTA).
@@
 Simultaneously, GPT-4 infers a reasonable instruction $\boldsymbol{p}^{\prime}$.
-When feeding the augmented instruction $\boldsymbol{p}_{i}^{\prime}$ and $\boldsymbol{p}_{j}^{\prime}$ rephrased by an LLM $\boldsymbol{G}_{3.5}$, the surrogate model $M$ extracts instruction-aware features of the AE $\boldsymbol{x}^{\prime}$ and $\boldsymbol{x}_{t}$, respectively.
-Finally, we minimize the $L_{2}$ distance between these two features to optimize the AE $\boldsymbol{x}^{\prime}$.
-Notably, the instruction $\boldsymbol{p}_{i/j}^{\prime}$ is updated on-the-fly during the optimization of the adversarial sample.
+Upon providing the augmented instruction $\boldsymbol{p}_{i}^{\prime}$ and $\boldsymbol{p}_{j}^{\prime}$ which are rephrased from $\boldsymbol{p}^{\prime}$ using GPT-4, the surrogate model $M$ extracts instruction-aware features of $\boldsymbol{x}_{t}$ and the AE $\boldsymbol{x}^{\prime}$, respectively.
+Finally, we minimize the $L_{2}$ distance between these two features to optimize $\boldsymbol{x}^{\prime}$.
 To overcome these challenges, this paper proposes an instruction-tuned targeted attack (InstructTA) for LVLMs with high attack transferability.
@@
 Our main contributions are outlined as follows:
-  • This paper explores the attack vector of LVLMs under a practical yet unexplored gray-box setting, i.e., the adversary only has the rights to access the visual encoder in LVLMs, excluding the LLM and instructions.
   • We propose an effective targeted attack, InstructTA.
-It cleverly leverages text-to-image models and LLMs to infer reasonable target images and instructions, respectively, and augments the transferability of adversarial samples with instruction-tuning paradigms.
+It cleverly leverages LLMs to deduce reasonable prompts for target texts, leading to extracting instruction-aware features for precise targeted attacks.
+  • We present the instruction-tuning paradigm in the targeted attacks to further augment the transferability of adversarial samples, benefiting from the rephrased instructions by GPT-4.
   • Extensive experiments demonstrate that InstructTA can achieve better targeted attack performance and transferability than intuitive baselines.
@@
 LVLMs.
-Benefiting from the powerful capabilities of LLMs in answering diverse and complex questions [31], LVLMs [46] have demonstrated remarkable performance on a variety of multi-modal tasks, e.g., image captioning, visual question-answering, natural language visual reasoning and visual dialog.
+Benefiting from the powerful capabilities of LLMs in answering diverse and complex questions [36], LVLMs [55] have demonstrated remarkable performance on a variety of multi-modal tasks, e.g., image captioning, visual question-answering, natural language visual reasoning, and visual dialog.
 To bridge the image-text modality gap, BLIP-2 [28] proposes the Q-Former to extract visual features into the frozen OPT [54] or FlanT5 [13].
-MiniGPT-4 [48] employs the identical pretrained ViT [16] and Q-Former as BLIP-2, but uses the more powerful Vicuna [10] to serve as the base LLM for text generation.
+MiniGPT-4 [57] employs the identical pre-trained ViT [18] and Q-Former as BLIP-2 but uses the more powerful Vicuna [12] to serve as the base LLM for text generation.
 InstructBLIP [14] introduces an instruction-aware Q-Former to extract the most task-relevant visual representations.
 LLaVA [32] bridges a pre-trained visual encoder with LLaMA, and fine-tunes them with question-image-answer paired data generated by GPT-4.
+Instead of projecting visual features onto the input embedding space of the LLM, CogVLM [46] introduces a visual expert module to synergistically incorporate image and linguistic features within the pre-trained large language model.
 Figure 2: The architecture of LVLM.
@@
 Although adversarial attacks are now being developed in a variety of modalities and tasks, the earliest and most extensive research occurred in the field of image classification [6, 44].
-In this case, the goal of adversarial attack is often to fool a well-trained visual classifier by crafted adversarial images.
+In this case, the goal of the adversarial attack is often to fool a well-trained visual classifier by crafted adversarial images.
 Based on their purpose, we divide adversarial attacks into targeted attacks and non-targeted attacks [1].
-Targeted attacks aim to confuse the model to recognize the adversarial example as a target label, whereas non-targeted attacks only ensure that the model make wrong predictions, not caring about the output error category.
-Moreover, according to the information of the attacked model exposed to the adversary, adversarial attacks can be grouped into white-box attack and black-box attack.
+Targeted attacks aim to confuse the model to recognize the adversarial example as a target label, whereas non-targeted attacks only ensure that the model makes wrong predictions, not caring about the output error category.
+Moreover, according to the information of the attacked model exposed to the adversary, adversarial attacks can be grouped into white-box attacks and black-box attacks.
 The white-box attacks can obtain the whole parameters and structure of the target model.
 Thus, they generally optimize adversarial perturbations with gradients of loss w.r.t inputs, e.g., FGSM [22], I-FGSM [26], C&W[7] and PGD [35].
-In contrast, only the input and output of the attacked model can be captured by black-box attack, which is more challenging and practical than white-box one.
+In contrast, only the input and output of the attacked model can be captured by black-box attacks, which is more challenging and practical than white-box one.
 One solution is to estimate the gradient or search for successful perturbations by directly querying the target model multiple times, e.g., ZOO [10] and NES attack [25].
@@
 AEs on LVLMs.
-Currently, there are some pioneering and contemporaneous works of targeted attacks on LVLMs.
+Currently, there are some pioneering works of targeted attacks on LVLMs.
 On one hand, some works aim to mislead the LVLMs to output specific words.
@@
 [42] also align the adversarial image and target images in the embedding space for LVLM jailbreak.
-Different from [46], we focus on the gray-box setting instead of the black-box setting, and further improve the attack transferability under the guidance of inferred and rephrased instructions.
-While [36] also accesses visual encoders in LVLMs for jailbreak, our work highlights the capability to implement generic targeted attacks even without knowing the target LVLM’s instructions.
+Different from [55], we further improve the attack transferability under the guidance of inferred and rephrased instructions.
+While [42] jailbreaks LVLMs with a generic prompt by perturbing the input image, our work highlights the capability to implement targeted attacks even without knowing the built-in prompts of the target LVLM.
 3 Method
 3.1 Problem Formulation
-LVLM. Given an image $\boldsymbol{x}$ and a text instruction $\boldsymbol{p}$ as inputs, the LVLM $F$ outputs an answer $\boldsymbol{y}$. Formally,
+LVLM. Given an image $\boldsymbol{x}$ and a text prompt $\boldsymbol{p}$ as inputs, the LVLM $F$ outputs an answer $\boldsymbol{y}$. Formally,
 $\boldsymbol{y}=F_{\theta}(\boldsymbol{x},\boldsymbol{p}),$ (1)
@@
 Typically, the LVLM [28, 57, 14, 32] projects the extracted visual features into the word embedding space and then feeds them to its underlying LLM.
-Hence, LVLM consists of a pre-trained visual encoder, a projector and a well-trained LLM, as shown in 2.
+Hence, LVLM consists of a pre-trained vision encoder, a connector and a well-trained LLM, as shown in Fig.
+2.
 Adversarial Objective.
@@
 s.t. $\|\boldsymbol{x}^{\prime}-\boldsymbol{x}\|_{\infty}\leq\epsilon,$
-where $\boldsymbol{x}$ is a given benign image sample, $\boldsymbol{x}^{\prime}$ represents the adversarial example of $\boldsymbol{x}$, and $\boldsymbol{y}_{t}$ denotes the predefined target text.
-$\epsilon$ is the budget of controlling adversarial perturbations, and $\boldsymbol{p}$ denotes the instructions used by the target LVLM ($\boldsymbol{p}$ is unknown to attackers; see § 3.2).
+where $\boldsymbol{x}$ is a given benign image, $\boldsymbol{x}^{\prime}$ represents the adversarial example of $\boldsymbol{x}$, and $\boldsymbol{y}_{t}$ denotes the predefined target text.
+$\epsilon$ is the budget of controlling adversarial perturbations, and $\boldsymbol{p}$ denotes the instruction used by the target LVLM ($\boldsymbol{p}$ is unknown to attackers; see § 3.2).
 $S(\cdot,\cdot)$ is a metric depicting the semantic similarity between two input sentences.
@@
 2 ensures that the response of the adversarial example shall retain the same meaning as the target text $y_{t}$.
-Algorithm 1 InstructTA
-0:  a benign image $\boldsymbol{x}$, a target text $\boldsymbol{y}_{t}$, a text-to-image generative model $h_{\xi}$, a surrogate model $M$, a VLP model with the visual encoder $f$ and text encoder $g$, ChatGPT $G_{3.5}$, GPT-4 $G_{4}$, custom prompts $\boldsymbol{p}_{g}$ and $\boldsymbol{p}_{r}$, step size $\eta$, perturbation budget $\epsilon$, attack iterations $T$
-1:  Initialize: $\boldsymbol{x}^{\prime}=\boldsymbol{x}$, $\boldsymbol{x}_{t}=h_{\xi}(\boldsymbol{y}_{t})$, $\boldsymbol{p^{\prime}}=G_{4}(\boldsymbol{p}_{g}\boldsymbol{y}_{t})$  // Eqn. (3)
-2:  for $iter=1...T$ do
-3:     $\boldsymbol{p}_{i}^{\prime}\leftarrow G_{3.5}(\boldsymbol{p}_{r}\boldsymbol{p}^{\prime})$
-4:     $\boldsymbol{p}_{j}^{\prime}\leftarrow G_{3.5}(\boldsymbol{p}_{r}\boldsymbol{p}^{\prime})$
-5:     $\boldsymbol{x}^{\prime}={\mathcal{S}}_{\epsilon}(\boldsymbol{x}^{\prime}-\eta\cdot\operatorname{sign}(\Delta_{\boldsymbol{x}^{\prime}}\mathcal{L}))$  // PGD
-6:  end for
-6:  adversarial example $\boldsymbol{x}^{\prime}$
 Figure 3: An example of rephrasing an instruction.
-Given the target response $\boldsymbol{y}_{t}$, GPT-4 infers an instruction, i.e., “Can you describe the scene depicted in the photograph?”.
-The real instruction assigned to this target text is “Explain the visual content of the image in great detail.”
+Given the target response $\boldsymbol{y}_{t}$, GPT-4 infers an instruction, i.e., “Can you describe what’s in the picture you’re looking at related to food?”.
+The real instruction assigned to this target text is “What are the essential components depicted in this image?”
 3.2 Targeted Attack with Instruction Tuning
@@
 Accordingly, attackers can locally prepare instructions $\boldsymbol{p}^{\prime}$ which are commonly used in the target domain (see below for details).
-We also assume that before attack, attackers can form a local surrogate model with the same visual encoder as that of the victim model.
-Inferring Instructions. We first locally infer the possible instructions $\boldsymbol{p}^{\prime}$ for the target text $\boldsymbol{y}_{t}$ by using GPT-4 [32]. Formally,
-$\boldsymbol{p^{\prime}}=G_{4}(\boldsymbol{p}_{g}\boldsymbol{y}_{t}),$ (3)
-where $G_{4}$ denotes the language-only GPT-4, and $\boldsymbol{p}^{\prime}$ represents an instruction generated by $G_{4}$.
-The above action denotes that we feed an instruction $\boldsymbol{p}_{g}$ concatenated with the target response text $\boldsymbol{y}_{t}$ to the GPT-4.
-In the experiment, we set $\boldsymbol{p}_{g}$ to a straightforward prompt “What is the question or prompt for this answer: ”.
+We also assume that before attack, attackers can form a local surrogate model with the same vision encoder as that of the victim model.
+Inferring Instructions.
+We first locally infer the possible instruction $\boldsymbol{p}^{\prime}$ for the target text $\boldsymbol{y}_{t}$ by using GPT-4 [37].
+We design the inference prompt $\boldsymbol{p}_{g}$ as follows.
+Inference prompt $\boldsymbol{p}_{g}$ What is the question or prompt for the following answer? You just provide a question or prompt. """ $\boldsymbol{y}_{t}$""
+According to the above prompt, we can derive a reasonable instruction $\boldsymbol{p}^{\prime}$ by feeding the target response text $\boldsymbol{y}_{t}$ to GPT-4.
 Forming Surrogate Models.
-Once the instructions have been inferred, the surrogate model uses a projector to transform the features from the visual encoder into instruction-aware features.
-Thus, the surrogate model comprises two modules: the visual encoder from the victim model and the pre-trained projector (e.g., Q-Former [25]).
-The projector maps global visual features into informative features tailored to given instructions.
+The surrogate model comprises two modules: the visual encoder sourced from the victim model and the pre-trained projector.
+The projector follows the visual encoder, as depicted in Fig.
+1.
+By providing an instruction as input, the projector is capable of mapping the existing global features into informative features that are specifically tailored to the given instruction.
+In this paper, Q-Former [28] is selected as the projector.
+Consequently, once the instructions have been inferred, the surrogate model employs the projector to convert the features obtained from the visual encoder into instruction-aware features.
 Generating Adversarial Examples.
@@
 Our tentative exploration shows that AEs generated by Eqn.
-4, although manifesting reasonable attackability, are often less transferable to the wide range of possible instructions in the wild.
-Thus, to improve the transferability and robustness of the generated adversarial examples, we use ChatGPT (i.e., GPT-3.5) to rephrase the inferred $\boldsymbol{p}^{\prime}$ as augmentation.
-The paraphrase prompt template fed into the ChatGPT is “Paraphrase this sentence: ”.
-This way, Eqn.
+3, although manifesting reasonable attackability, are often less transferable to the wide range of possible instructions in the wild.
+Thus, to improve the transferability and robustness of the generated adversarial examples, we use GPT-4 to rephrase the inferred $\boldsymbol{p}^{\prime}$ as augmentation.
+The prompt $\boldsymbol{p}_{r}$ for paraphrasing is formulated as follows.
+Rephrasing prompt $\boldsymbol{p}_{r}$ Rephrase the following sentence without changing its original meaning.
+Please give $n-1$ examples of paraphrasing that look different and begin the statement with "-".
+""" $\boldsymbol{p}^{\prime}$""
+where $n-1$ represents the number of sentences rephrased by GPT-4.
+Together with the original inferred instruction $\boldsymbol{p}^{\prime}$, a total of $n$ instructions are collected in a set $Q$.
+We set $n$ to 10 by default.
+Benefiting from the augmented instructions, Eqn.
 3 can be rewritten as
-$\min_{\boldsymbol{x}^{\prime}}\mathcal{J}_{ins}=\mathbb{E}_{\boldsymbol{p}_{i}^{\prime},\boldsymbol{p}_{j}^{\prime}}\|M(\boldsymbol{x}^{\prime}$ $,\boldsymbol{p}_{i}^{\prime})-M(h_{\xi}(\boldsymbol{y}_{t}),\boldsymbol{p}_{j}^{\prime})\|_{2}^{2},$ (5)
+$\min_{\boldsymbol{x}^{\prime}}\mathcal{J}_{ins}=\mathbb{E}_{\boldsymbol{p}_{i}^{\prime},\boldsymbol{p}_{j}^{\prime}\in Q}\|M(\boldsymbol{x}^{\prime}$ $,\boldsymbol{p}_{i}^{\prime})-M(h_{\xi}(\boldsymbol{y}_{t}),\boldsymbol{p}_{j}^{\prime})\|_{2}^{2},$ (4)
 […]
-where $\boldsymbol{p}_{i}^{\prime}$ and $\boldsymbol{p}_{j}^{\prime}$ are augmented instructions from $\boldsymbol{p}^{\prime}$.
-Both of them are generated by ChatGPT, i.e., $\boldsymbol{p}_{i}^{\prime}=G_{3.5}(\boldsymbol{p}_{r}\boldsymbol{p}^{\prime})$, $\boldsymbol{p}_{j}^{\prime}=G_{3.5}(\boldsymbol{p}_{r}\boldsymbol{p}^{\prime})$, where $G_{3.5}$ represents ChatGPT and $\boldsymbol{p}_{r}$ denotes the aforementioned rephrasing template.
+where $\boldsymbol{p}_{i}^{\prime}$ and $\boldsymbol{p}_{j}^{\prime}$ are randomly selected from the set $Q$.
+We utilize the rephrasing prompt $\boldsymbol{p}_{r}$ to construct the instruction set $Q$ via GPT-4.
 As illustrated in Fig.
-3, ChatGPT is generally effective to rephrase a seed instruction $\boldsymbol{p}^{\prime}$ into a set of semantically-correlated instructions.
-Table 1: Targeted attacks against victim LVLMs.
-A total of 1,000 clean images denoted as $\boldsymbol{x}$ are sampled from the ImageNet-1K [13] validation set.
-For each of these benign images, a targeted text $\boldsymbol{y}_{t}$ and a prompt $\boldsymbol{p}$ are randomly assigned from LLaVA-Instruct-150K [27].
-Our analysis involves the computation of the CLIP score (normal-$\uparrow$) [34, 46] between the target texts and the generated responses of AEs on attacked LVLMs, which was performed using several CLIP text encoders and their ensemble.
-We also report the No.
-of successful attacks within 1,000 adversarial samples, determined by GPT-4.
-CLIP and EVA-CLIP in parentheses indicate that they are used as surrogate models for transfer attack.
-VLVM Attack method Text encoder (pretrained) for evaluation The No.
-of successes RN50 RN101 ViT-B/16 ViT-B/32 ViT-L/14 Ensemble BLIP-2 Clean image 0.296 0.427 0.342 0.354 0.171 0.318 4 MF-it (CLIP) [46] 0.292 0.424 0.339 0.353 0.168 0.315 2 MF-ii (CLIP) [46] 0.293 0.426 0.340 0.353 0.168 0.316 2 MF-it (EVA-CLIP) 0.574 0.667 0.616 0.630 0.481 0.594 519 MF-ii (EVA-CLIP) 0.573 0.661 0.616 0.627 0.484 0.592 514 InstructTA-MF-it (ours) 0.604 0.691 0.644 0.656 0.515 0.622 652 InstructTA-MF-ii (ours) 0.599 0.682 0.639 0.649 0.512 0.616 619 InstructBLIP Clean image 0.392 0.557 0.420 0.421 0.312 0.420 0 MF-it (CLIP) [46] 0.394 0.558 0.423 0.422 0.313 0.422 0 MF-ii (CLIP) [46] 0.395 0.557 0.424 0.423 0.314 0.422 0 MF-it (EVA-CLIP) 0.773 0.822 0.791 0.796 0.740 0.785 292 MF-ii (EVA-CLIP) 0.761 0.813 0.780 0.783 0.727 0.773 259 InstructTA-MF-it (ours) 0.810 0.849 0.826 0.830 0.784 0.820 383 InstructTA-MF-ii (ours) 0.789 0.833 0.805 0.809 0.758 0.799 328 MiniGPT-4 Clean image 0.345 0.495 0.367 0.374 0.253 0.367 0 MF-it [46] 0.341 0.493 0.364 0.369 0.251 0.364 0 MF-ii [46] 0.340 0.492 0.363 0.369 0.249 0.362 2 MF-it (EVA-CLIP) 0.650 0.726 0.664 0.676 0.593 0.662 259 MF-ii (EVA-CLIP) 0.649 0.723 0.665 0.670 0.592 0.660 247 InstructTA-MF-it (ours) 0.692 0.756 0.701 0.712 0.638 0.700 363 InstructTA-MF-ii (ours) 0.680 0.748 0.692 0.701 0.627 0.690 325
+3, GPT-4 is generally effective to rephrase a seed instruction $\boldsymbol{p}^{\prime}$ into a set of semantically-correlated instructions.
 3.3 Dual Targeted Attack & Optimization
-With the aforementioned steps, we form following dual targeted attack framework:
+With the aforementioned steps, we form the following dual targeted attack framework:
 $\arg\min_{\boldsymbol{x}^{\prime}}\mathcal{L}=\mathcal{J}_{ins}-\mathcal{J}_{mf},{\quad\text{s.t.}~{}}\|\boldsymbol{x}^{\prime}&-\boldsymbol{x}\|_{\infty}\leq\epsilon\end{aligned},$ (5)
-where $\mathcal{J}_{mf}$ is the objective of MF-it or MF-ii, following [46].
-MF-it and MF-ii achieve transfer targeted attack via matching the embeddings of the adversarial example and the target response.
+where $\mathcal{J}_{mf}$ is the objective of MF-it, following [55].
+MF-it achieves transfer targeted attack via matching the embeddings of the adversarial example and the target response.
 For MF-it, its objective is defined as
@@
 where $f(\cdot)$ and $g(\cdot)$ are the paired visual encoder and text encoder of an align-based VLP model (e.g., CLIP [40], ALBEF [29], Open-CLIP [11], and EVA-CLIP [19]), respectively.
-For MF-ii, we have
-[…]
-Notably, Eqn. 6 with MF-it is marked InstructTA-MF-it, and Eqn. 6 with MF-ii is marked InstructTA-MF-ii.
-When optimizing adversarial samples following Eqn. 6, this paper adopts PGD to update $\boldsymbol{x}^{\prime}$ with $T$ iterations as below,
+When optimizing adversarial samples following Eqn. 5, this paper adopts PGD to update $\boldsymbol{x}^{\prime}$ with $T$ iterations as below,
 $\boldsymbol{x}_{T}^{\prime}={\mathcal{S}}_{\epsilon}(\boldsymbol{x}_{T-1}^{\prime}-\eta\cdot\operatorname{sign}(\Delta_{\boldsymbol{x}_{T-1}^{\prime}}\mathcal{L})),\quad\boldsymbol{x}^{\prime}_{0}=\boldsymbol{x},$ (7)
@@
 Initially, once decided an adversarial target text $\boldsymbol{y}_{t}$, we use the text-to-image model $h_{\xi}$ and GPT-4 to “reverse” a target image $\boldsymbol{x}_{t}$ and produce the reference instruction $\boldsymbol{p}^{\prime}$, respectively.
-In each iteration, we augment the inferred instruction $\boldsymbol{p}^{\prime}$ with ChatGPT firstly.
+Furthermore, we augment the inferred instruction $\boldsymbol{p}^{\prime}$ with GPT-4 to form the instruction set $Q$.
+In each iteration, we randomly select instructions $\boldsymbol{p}_{i}^{\prime}$ and $\boldsymbol{p}_{j}^{\prime}$ from $Q$.
 Then, we update the adversarial example with PGD over Eqn.
-6.
+5.
 Until the end of $T$ step iterations, the final adversarial sample is obtained and used for attacking remote LVLMs.
+Algorithm 1 InstructTA
+0:  a benign image $\boldsymbol{x}$, a target text $\boldsymbol{y}_{t}$, a text-to-image generative model $h_{\xi}$, a surrogate model $M$, a VLP model with the visual encoder $f$ and text encoder $g$, GPT-4 $G_{4}$, custom prompts $\boldsymbol{p}_{g}$ and $\boldsymbol{p}_{r}$, step size $\eta$, perturbation budget $\epsilon$, attack iterations $T$
+1:  Initialize: $\boldsymbol{x}^{\prime}=\boldsymbol{x}$, $\boldsymbol{x}_{t}=h_{\xi}(\boldsymbol{y}_{t})$,   // generate a target image $\boldsymbol{p^{\prime}}=G_{4}(\boldsymbol{p}_{g}\boldsymbol{y}_{t})$,  //infer an instruction $Q\leftarrow G_{4}(\boldsymbol{p}_{r}\boldsymbol{p^{\prime}})$  // build an instruction set $Q$ by rephrasing $\boldsymbol{p^{\prime}}$
+2:  for $iter=1...T$ do
+3:     $\boldsymbol{p}_{i}^{\prime}\sim Q$, $\boldsymbol{p}_{j}^{\prime}\sim Q$  // randomly select instructions
+4:     $\mathcal{L}=\mathcal{J}_{ins}-\mathcal{J}_{mf}$  // Eqn. 5
+5:     $\boldsymbol{x}^{\prime}={\mathcal{S}}_{\epsilon}(\boldsymbol{x}^{\prime}-\eta\cdot\operatorname{sign}(\Delta_{\boldsymbol{x}^{\prime}}\mathcal{L}))$  // PGD
+6:  end for
+6:  adversarial example $\boldsymbol{x}^{\prime}$
 4 Experiments
@@
 The ImageNet-1K [15] comprises 3 subsets, including training, validation and test sets.
-We randomly select 1,000 images from the validation set to serve as the benign samples ($\boldsymbol{x}$) for the attack.
-Correspondingly, we sample 1,000 instruction-answer pairs from LLaVA-Instruct-150K [27] as tested instructions and target texts, respectively, since LLaVA-Instruct-150K is a set of multimodal instruction-following data generated by GPT-4.
+We randomly select 1,000 images from the validation set to serve as the benign samples ($\boldsymbol{x}$) for the targeted attack.
+To establish the instruction-answer pairs as tested instructions and target texts, we adapt 11 instructions (see the supplementary material) from LLaVA-Instruct-150K [32] and employ GPT-4 to generate the corresponding answers.
+Specifically, we pick 1,000 images from the validation set of COCO [30], and for each image, we randomly select an instruction from these 11 instructions and utilize the GPT-4V [38] API to output the response.
 In addition, we use Stable Diffusion ($h_{\xi}$) to generate target images from the target texts, following [55].
@@
 Following [55], we adopt CLIP-score ($\uparrow$) to evaluate the targeted attack performance, which calculates the semantic similarity between the generated response of victim LVLM and the target text using CLIP text encoders.
-Additionally, we record the number of successful attacks by using GPT-4 to determine whether two texts have the same meaning.
-The prompt in this operation is as follows: Determine whether these two texts describe the same objects or things, you only need to answer yes or no: 1.
-$\boldsymbol{y}^{\prime}$ 2.
-$\boldsymbol{y}_{t}$ where $\boldsymbol{y}^{\prime}$ denotes the output response of the adversarial example on victim model, and $\boldsymbol{y}_{t}$ is the target response.
+Additionally, we record the number of successful attacks (NoS) and attack success rate (ASR) by using GPT-4 to determine whether two texts have the same meaning.
+The prompt in this operation is as follows:
+Attack success prompt Determine whether these two texts describe the same objects or things, you only need to answer yes or no: 1. $\boldsymbol{y}^{\prime}$ 2. $\boldsymbol{y}_{t}$
+where $\boldsymbol{y}^{\prime}$ denotes the output response of the adversarial example on the victim model, and $\boldsymbol{y}_{t}$ is the target response.
 Baselines.
-This paper provides evaluations on three victim LVLMs, including BLIP-2 [25], InstructBLIP [12] and MiniGPT-4 [48].
-In detail, we select flan-t5-xxl [^2] https://huggingface.co/google/flan-t5-xxl, vicuna-7b-v1.1333https://huggingface.co/lmsys/vicuna-7b-v1.1 and vicuna-v0-13b [^4] https://huggingface.co/Vision-CAIR/vicuna as LLMs for BLIP-2, InstructBLIP and MiniGPT-4, respectively.
-Their visual encoders are all ViT-G/14 [42] borrowed from EVA-CLIP [17].
-To evaluate the targeted attack performance, InstructTA is compared with existing baselines, i.e., MF-it and MF-ii.
-They can choose CLIP (ViT-B/32) [34] or the EVA-CLIP (ViT-G/14) as surrogate models for targeted attacks.
+This paper provides evaluations on five victim LVLMs, including BLIP-2 [28], InstructBLIP [14], MiniGPT-4 [57], LLaVA [32], and CogVLM [46].
+In detail, we select flan-t5-xxl [^2] https://huggingface.co/google/flan-t5-xxl, vicuna-7b-v1.1333https://huggingface.co/lmsys/vicuna-7b-v1.1, vicuna-v0-13b [^4] https://huggingface.co/Vision-CAIR/vicuna, LLaMA-2-13B [^5] https://huggingface.co/liuhaotian/llava-llama-2-13b-chat-lightning-preview and vicuna-7b-v1.5666https://huggingface.co/lmsys/vicuna-7b-v1.5 as LLMs for BLIP-2, InstructBLIP, MiniGPT-4, LLaVA and CogVLM [^7] https://huggingface.co/THUDM/cogvlm-base-224-hf, respectively.
+For BLIP-2, InstructBLIP and MiniGPT-4, their visual encoders are all ViT-G/14 [51] borrowed from EVA-CLIP [19].
+The vision encoders for LLaVA and CogVLM are built on ViT-L/14 [40] and EVA2-CLIP-E [43], respectively.
+To evaluate the targeted attack performance, InstructTA is compared with existing baselines, i.e., MF-tt, MF-it and MF-ii [55].
+MF-tt is a query-based black-box attack method that only accesses the input and output of the target model.
 Implementation Details.
 All images are resized to $224\times 224$ before feeding in substitute models and LVLMs.
-For InstructTA, we choose the visual encoder and Q-Former from InstructBLIP as the surrogate model.
+MF-it and MF-ii choose the same vision encoders as the victim LVLMs as surrogate models for targeted attacks.
+For InstructTA, we choose the vision encoder (ViT-G/14) and Q-Former from InstructBLIP as the surrogate model.
+In other words, ViT-G/14 and Q-Former are the visual encoder and projector in Fig.
+1, respectively.
+For BLIP-2, InstructBLIP and MiniGPT-4, our surrogate models of InstructTA share the same vision encoders as the target models, regardless of the instruction-aware surrogate model or MF-it used in the dual targeted attack.
+For LLaVA and CogVLM, only MF-it in InstructTA (Eqn.
+5) keeps the same vision encoder as the victim model.
 For PGD, $\eta$ and $T$ are set to 1 and 100, respectively.
 The maximum perturbation magnitude $\epsilon$ is set to $8$.
+The settings of MF-tt, MF-it and MF-ii are all consistent with the original literature [55].
 We implement all the evaluations using Pytorch 2.1.0 and run them on a single NVIDIA RTX A6000 GPU (48G).
+Table 1: Targeted attacks against victim LVLMs.
+A total of 1,000 clean images denoted as $\boldsymbol{x}$ are sampled from the ImageNet-1K [15] validation set.
+For each of these benign images, a targeted text $\boldsymbol{y}_{t}$ and a prompt $\boldsymbol{p}$ are randomly assigned from our generated instruction-answer paired set (see § 4.1).
+Our analysis involves the computation of the CLIP-score ($\uparrow$) [40, 55] between the target texts and the generated responses of AEs on attacked LVLMs, which was performed using several CLIP text encoders and their ensemble.
+We also report the No.
+of successful attacks (NoS) within 1,000 adversarial samples, determined by GPT-4.
+LVLM Attack method Text encoder (pretrained) for evaluation NoS ($\uparrow$) RN50 RN101 ViT-B/16 ViT-B/32 ViT-L/14 Ensemble BLIP-2 Clean image 0.281 0.395 0.323 0.333 0.184 0.303 1 MF-tt [55] 0.284 0.400 0.326 0.337 0.185 0.306 0 MF-it [55] 0.533 0.634 0.575 0.581 0.469 0.559 401 MF-ii [55] 0.536 0.634 0.575 0.583 0.471 0.560 414 InstructTA 0.562 0.660 0.604 0.611 0.502 0.588 519 InstructBLIP Clean image 0.329 0.450 0.358 0.361 0.246 0.349 1 MF-tt [55] 0.332 0.452 0.361 0.363 0.249 0.351 1 MF-it [55] 0.650 0.714 0.675 0.676 0.608 0.665 233 MF-ii [55] 0.638 0.707 0.666 0.667 0.594 0.654 223 InstructTA 0.683 0.740 0.709 0.708 0.645 0.697 300 MiniGPT-4 Clean image 0.324 0.448 0.348 0.351 0.244 0.343 2 MF-tt [55] 0.332 0.455 0.353 0.356 0.243 0.348 0 MF-it [55] 0.593 0.672 0.611 0.618 0.538 0.606 121 MF-ii [55] 0.602 0.682 0.621 0.623 0.548 0.615 128 InstructTA 0.633 0.701 0.650 0.654 0.578 0.643 164 LLaVA Clean image 0.325 0.465 0.374 0.383 0.216 0.352 1 MF-tt [55] 0.329 0.466 0.376 0.386 0.217 0.355 1 MF-it [55] 0.368 0.502 0.418 0.424 0.268 0.396 32 MF-ii [55] 0.360 0.496 0.407 0.415 0.256 0.387 30 InstructTA 0.395 0.524 0.445 0.449 0.300 0.423 36 CogVLM Clean image 0.301 0.425 0.319 0.334 0.208 0.318 4 MF-tt [55] 0.307 0.432 0.327 0.342 0.212 0.324 8 MF-it [55] 0.493 0.597 0.506 0.513 0.417 0.505 71 MF-ii [55] 0.499 0.600 0.516 0.521 0.423 0.512 63 InstructTA 0.528 0.624 0.539 0.545 0.452 0.537 103
 (a) The target text and image
 (b) Clean image
-(c) MF-it (CLIP)
-(d) MF-ii (CLIP)
-(e) MF-it (EVA-CLIP)
-(f) MF-ii (EVA-CLIP)
-(g) InstructTA-MF-it
-(h) InstructTA-MF-ii
-Figure 4: Visualization examples of various targeted attack methods on InstructBLIP. “Explain the visual content of the image in great detail.” is a real instruction.
+(c) MF-tt
+(d) MF-it
+(e) MF-ii
+(f) InstructTA
+Figure 4: Visualization examples of various targeted attack methods on InstructBLIP. “What are the essential components depicted in this image?" is a real instruction.
 4.2 Results
@@
 As illustrated in Table 1, we transfer adversarial samples generated by different targeted attack methods to victim LVLMs and mislead them to produce target responses.
-For each adversarial example, when attacking a LVLM, the target text and the assigned instruction may be different from other adversarial examples, since we randomly selected 1000 pairs of target texts and instructions from LLaVA-Instruct-150K [27].
+For each adversarial example, when attacking an LVLM, the target text and the assigned instruction may be different from other adversarial examples, since we randomly selected 1000 pairs of target texts and instructions from our generated instruction-answer paired set (see § 4.1).
 Subsequently, we evaluate the similarity between the generated response $\boldsymbol{y}^{\prime}$ of the victim LVLM and the targeted text $\boldsymbol{y}_{t}$ by utilizing diverse forms of CLIP text encoders, where $\boldsymbol{y}^{\prime}=F(\boldsymbol{x}^{\prime},\boldsymbol{p})$.
-In addition, we count the number of successful targeted adversarial samples, as judged by GPT-4.
+In addition, we tabulate the count of successful targeted adversarial samples (NoS), as evaluated by GPT-4.
 Table 1 reports the evaluation results of different targeted attack methods on BLIP-2, InstructBLIP and MiniGPT-4.
-We compare InstructTA with MF-it and MF-ii, which are the two recent attacks on LVLMs.
-Overall, we observe that InstructTA consistently outperforms MF-it and MF-ii in terms of CLIP score and the number of successful attacks.
-In fact, it can be seen that MF-it and MF-ii with CLIP have almost the same results as the clean samples, indicating no targeted attack effect.
-We believe this is reasonable, because ViT-B/32 in CLIP shares a different visual encoder than that of the target LVLM.
-In contrast, MF-it (EVA-CLIP), MF-ii (EVA-CLIP) and our methods all notably outperform the targeted attack results on clean samples because they share the same visual encoder as the victim model.
-Overall, we interpret these results justify the necessity of using the same visual encoder as the victim model.
-More importantly, both InstructTA-MF-it and InstructTA-MF-ii are distinctly better than MF-it and MF-ii in terms of CLIP-score and the number of successful attacks.
-These phenomena demonstrate that our method can effectively improve the transferability of the targeted attack on LVLMs.
-One reason is that we enhance the robustness of AEs to varying instructions.
-Another reason is that our dual attacks (InstructTA-MF-it/InstructTA-MF-ii) extracts informative features containing global and instruction-aware knowledges, benefiting to precise targeted attacks.
+We compare InstructTA with MF-tt, MF-it and MF-ii, which are the three recent attacks on LVLMs.
+Overall, we observe that InstructTA consistently outperforms MF-tt, MF-it and MF-ii in terms of CLIP score and NoS.
+In fact, it can be seen that MF-tt has almost the same results as the clean samples, indicating no promising targeted attack effect.
+In contrast, MF-it, MF-ii, and our method exhibit significant improvements over the clean samples on targeted attack results.
+This observation indicates the effectiveness of targeted attack methods against victim LVLMs.
+Significantly, InstructTA surpasses MF-it and MF-ii in terms of CLIP-score and NoS, underscoring its superiority.
+These observations highlight the capacity of our method to enhance the transferability of targeted attacks on LVLMs.
+This is attributable to two factors.
+Firstly, we enhance the resilience of adversarial examples to diverse instructions, thereby bolstering their robustness.
+Secondly, our dual attack scheme extracts informative features that encompass both global and instruction-aware knowledge, thus facilitating precise targeted attacks.
 Table 2: Ablation study of the proposed method on BLIP-2.
-CLIP score [46] is computed by ensemble CLIP text encoders.
-Attack success rate is the ratio between the No.
+CLIP-score [55] is computed by ensemble CLIP text encoders.
+Attack success rate (ASR) is the ratio between the No.
 of attack successes and the total No.
 of samples.
-Attack method CLIP score ($\uparrow$) Attack success rate ($\uparrow$) InstructTA-MF-it 0.622 0.652 InstructTA-MF-ii 0.616 0.619 InstructTA 0.606 0.624 InstructTA-woG 0.571 0.448
+         Attack method          CLIP-score ($\uparrow$)          ASR ($\uparrow$)          InstructTA          0.588          0.519          InstructTA-woMF          0.574          0.472          InstructTA-woMFG          0.521          0.324
+Generalization of InstructTA.
+We argue that InstructTA exhibits a degree of generalization in attacking LVLMs, even when the vision encoder of the instruction-aware surrogate model differs from that of the target LVLM.
+As evidenced in Table 1, the surrogate model employed is ViT-G/14 from EVA-CLIP [19], distinct from the ViT-G/14 and EVA02-bigE-14 utilized in LLaVA and CogVLM respectively.
+Despite this discrepancy, we observe that InstructTA outperforms previous methods in terms of targeted attack effectiveness.
+Hence, InstructTA can enhance its attack transferability beyond MF-it without necessitating the maintenance of an identical visual encoder to the victim LVLM.
 Visualization Examples.
 Fig.
-3 has displayed a sample inferred instruction and its sibling instructions rephrased by ChatGPT.
-It can be seen that the instructions inferred by GPT-4 are similar to the real instruction.
+3 has displayed a sample inferred instruction and its sibling instructions rephrased by GPT-4.
+It can be seen that the instruction inferred by GPT-4 is similar to the real instruction.
 The rephrased instructions are even more similar to the original sentence.
@@
 4.
-Overall, it can be observed that the generated responses of our methods are closer to the target texts.
+Overall, it can be observed that the generated response of our methods is closer to the target text.
 Table 3: Targeted attack performance for multiple perturbation budget $\epsilon$.
-Budget $\epsilon$ Text encoder (pretrained) for evaluation Attack success rate RN50 RN101 ViT-B/16 ViT-B/32 ViT-L/14 Ensemble 2 0.437 0.547 0.480 0.492 0.330 0.457 0.154 4 0.560 0.651 0.602 0.612 0.470 0.579 0.450 8 0.604 0.691 0.644 0.656 0.515 0.622 0.652 16 0.617 0.702 0.657 0.669 0.527 0.634 0.746 64 0.617 0.704 0.658 0.671 0.529 0.636 0.740
-Figure 5: To explore the impact of varying $\epsilon$ values within the InstructTA-MF-it, we conducted experiments aiming to achieve different levels of perturbed images on BLIP-2, i.e., referred to as the AE $\boldsymbol{x}^{\prime}$.
+Budget $\epsilon$ Text encoder (pretrained) for evaluation ASR ($\uparrow$) RN50 RN101 ViT-B/16 ViT-B/32 ViT-L/14 Ensemble 2 0.406 0.516 0.451 0.458 0.326 0.431 0.127 4 0.510 0.613 0.552 0.558 0.445 0.535 0.349 8 0.562 0.660 0.604 0.611 0.502 0.588 0.519 16 0.570 0.669 0.610 0.617 0.509 0.595 0.605 64 0.575 0.673 0.616 0.623 0.516 0.600 0.604
+Figure 5: To explore the impact of varying $\epsilon$ values within the InstructTA, we conducted experiments aiming to achieve different levels of perturbed images on BLIP-2, i.e., referred to as the AE $\boldsymbol{x}^{\prime}$.
 Our findings indicate a degradation in the visual quality of $\boldsymbol{x}^{\prime}$, as quantified by the LPIPS [53] distance between the original image $\boldsymbol{x}$ and the adversarial image $\boldsymbol{x}^{\prime}$.
@@
 4.3 Ablation Study
-We also launched ablation study to explore the contribution of major components in the proposed method.
-At this step, we conducted experiments on BLIP-2 with two key component ablated sequentially: 1) removing MF-it and MF-ii (denoted as InstructTA), and 2) launching optimization without CLIP embedding matching nor ChatGPT for paraphrase (in accordance with Eqn.
-4; denoted as InstructTA-woG).
+We also launched an ablation study to explore the contribution of major components in the proposed method.
+At this step, we conducted experiments on BLIP-2 with two key components ablated sequentially: 1) removing MF-it (denoted as InstructTA-woMF), and 2) launching optimization without MF-it nor GPT-4 for paraphrases (in accordance with Eqn.
+3; denoted as InstructTA-woMFG).
 The results are reported in Table 2.
 It is observed that the performances degrade when one or more parts are removed.
-In particular, the InstructTA-woG setting exhibits a success rate downgrading of over 31% in comparison to the InstructTA-MF-it setting.
+In particular, the InstructTA-woMFG setting exhibits a success rate downgrading of over 19% in comparison to the complete InstructTA framework.
 Overall, we interpret the ablation study results as reasonable, illustrating that all the constituent components have been thoughtfully devised and exhibit congruity, culminating in an efficacious solution for the targeted attack.
@@
 Effect of $\epsilon$.
-To explore the effect of the perturbation magnitude $\epsilon$ on targeted attack performance, we study the attack results of InstructTA-MF-it toward BLIP-2 under various $\epsilon$.
+To explore the effect of the perturbation magnitude $\epsilon$ on targeted attack performance, we study the attack results of InstructTA toward BLIP-2 under various $\epsilon$.
 The results are reported in Table 3.
-It is evident that with $\epsilon$ increasing, both CLIP scores and attack success rates rise, illustrating stronger attack performance.
+It is evident that with $\epsilon$ increasing, both CLIP-scores and attack success rates rise, illustrating stronger attack performance.
 Nevertheless, it is seen that the attack performance becomes gradually saturated when $\epsilon$ is greater than $16$.
@@
 Thus, we suggest to explore a proper $\epsilon$ for the targeted attack, e.g., $\epsilon=8$, such that the attack performance can be effectively improved while the perturbation is imperceptible.
-Table 4: Average CLIP score ($\uparrow$) between different versions of instructions. Each type of instruction have 1,000 samples.
-Instruction Rephrased instruction $\boldsymbol{p}_{i/j}^{\prime}$ Real instruction $\boldsymbol{p}$ Inferred instruction $\boldsymbol{p}^{\prime}$ 0.942 0.883 Real instruction $\boldsymbol{p}$ 0.882 1.000
+Effect of count $n$.
+In rephrasing prompt $\boldsymbol{p}_{r}$, we use $n-1$ to control the number of instructions rephrased by GPT-4.
+With the original instruction, we totally collect $n$ instructions to participate in the tuning for our targeted attack.
+To explore the effect of $n$ on targeted attack performance, we perform InstructTA on BLIP-2 with varying $n$.
+The results are illustrated in Table 4.
+We can observe that CLIP scores are roughly the same for different $n$.
+It is worth noting that as n increases, the attack success rate (ASR) also increases.
+Therefore, increasing the number of paraphrased instructions can further improve ASR.
+Table 4: Targeted attack performance for different $n$.
+Parameter $n$ Text encoder (pretrained) for evaluation ASR ($\uparrow$) RN50 RN101 ViT-B/16 ViT-B/32 ViT-L/14 Ensemble 1 0.560 0.658 0.600 0.606 0.498 0.584 0.504 5 0.560 0.657 0.600 0.606 0.498 0.584 0.504 10 0.562 0.660 0.604 0.611 0.502 0.588 0.519 50 0.562 0.659 0.601 0.607 0.499 0.586 0.529
+Table 5: Average CLIP-score ($\uparrow$) between different versions of instructions. Each type of instruction has 1,000 samples.
+Instruction Rephrased instruction $\boldsymbol{p}_{i/j}^{\prime}$ Real instruction $\boldsymbol{p}$ Inferred instruction $\boldsymbol{p}^{\prime}$ 0.943 0.829 Real instruction $\boldsymbol{p}$ 0.828 1.000
 Assessment of Generated Instructions.
-To quantitatively evaluate the usability of the generated instructions, we collect a paraphrased instruction for each inferred instruction used in main experiment and provide average CLIP score values between them, as reported in Table 4.
+To quantitatively evaluate the usability of the generated instructions, we collect a paraphrased instruction for each inferred instruction used in the main experiment and provide average CLIP-score values between them, as reported in Table 5.
 It is evident that there exists a high similarity score between the rephrased and the inferred instructions; moreover, the rephrased and the inferred instructions both show high similarity with the real instructions.
@@
 For instance, we can leverage the adversarial training framework [35] to train LVLMs with adversarial examples.
-We can also explore the possibility of incorporating adversarial training into the pre-training stage of LVLMs or their visual encoders.
+We can also explore the possibility of incorporating adversarial training into the pre-training stage of LVLMs or their vision encoders.
 Moreover, we can also explore detecting adversarial examples by leveraging the discrepancy between the instruction-aware features of the adversarial examples and the original images.
-Supplementary Material
-6 Attack Cases
-Fig.
-6 shows an example to attack BLIP-2 for different methods.
-It can be seen from the figure that CLIP-based MF-it and MF-ii shows no targeted attack effect since they have the similar results with the clean sample.
-Interestingly, half of the response (i.e., “a man and a woman”) of the adversarial sample generated by MF-it (EVA-CLIP) matches the original image and half (i.e., “elephants in a zoo”) matches the target image.
-MF-ii (EVA-CLIP), InstructTA-MF-it and InstructTA-MF-ii demonstrate the targeted attack effect.
-Importantly, the description of our InstructTA-MF-ii best matches the target image.
-7 Results with Complex Reasoning Questions
-Inferring from images can demonstrate the model’s deep understanding of the image semantics.
-The box below shows some reasoning instructions from complex_reasoning_77k in LLaVA-Instruct-150K [27].
-1.
-How can pedestrians ensure their safety when navigating this busy city scene at night?
-2.
-What might the cat be curious about on the floor?
-3.
-What suggestions can be made for individuals wishing to serve these dough balls in an aesthetically pleasing manner?
-4.
-What might be the man’s intention in the kitchen?
-5.
-How does the position of the players on the field affect their performance?
-6.
-What unique feature can be observed about the desserts in the scene?
-7.
-What type of event might be taking place in this scene?
-8.
-What could be a possible reason for the dog and cat’s interaction?
-9.
-In terms of architectural features, why might this clock tower be studied or admired?
-10.
-What tactic might the man be employing during the tennis match?
-11.
-What could be the significance or purpose of a single pink rose placed in a glass with water?
-12.
-What can you infer about the laptop owner’s personal style or interests?
-13.
-How is the truck being used during this event, and what challenges might it face in this context?
-14.
-Why might the vandalized stop sign on the school bus cause concern?
-…
-To further evaluate the effectiveness of the proposed methods, we provide transfer attack results for complex reasoning instructions, as portrayed in Table 6.
-All methods, even clean samples, have obvious targeted attack performance.
-This may be because some of the selected reasoning instructions can be answered without the need for image information, e.g., “Why might the vandalized stop sign on the school bus cause concern?”.
-Importantly, our methods have better transfer attack performances in terms of both CLIP score and the No.
-of successes.
-It implies that InstructTA still performs well even when faced with complex inference prompts.
-8 Victim Models
-Table 5 shows the architectures of the target LVLMs.
-Table 5: The architectures of the victim LVLMs.
-LVLM Visual Encoder Projector LLM BLIP-2 ViT-G/14 from EVA-CLIP Q-Former & Single linear layer Flan-T5-XXL InstructBLIP ViT-G/14 from EVA-CLIP Q-Former & Single linear layer Vicuna-7b-v1.1 MiniGPT-4 ViT-G/14 from EVA-CLIP Q-Former & Single linear layer Vicuna-v0-13b
-Figure 6: An instance of different targeted attack methods on BLIP-2.
-The real instruction is “Can you describe the main features of this image for me?”.
-The target text is “The image depicts a large elephant and a smaller elephant standing in a grassy area next to a stone wall, likely within a zoo enclosure.
-The two elephants appear to be gathered near a flock of black birds.
-The birds can be seen scattered around the enclosure, with some standing near the elephants and others in the background.
-The scene captures the interaction between the elephants and the birds, creating a serene and captivating moment.”
-Table 6: Targeted attacks against victim BLIP-2 with reasoning instructions.
-We randomly select a total of 1,000 clean images from the ImageNet-1K [13] validation set.
-For each of these benign sample, a target text and a prompt are randomly assigned from complex_reasoning_77k in LLaVA-Instruct-150K [27].
-Our metrics involves the computation of the CLIP score (normal-$\uparrow$) [34, 46] between the target texts and the generated responses of AEs on the attacked LVLM, which was performed using several CLIP text encoders and their ensemble.
-We also report the No.
-of successful attacks within 1,000 adversarial samples, determined by GPT-4.
-CLIP and EVA-CLIP in parentheses indicate that they are used as surrogate models for transfer attack.
-VLVM Attack method Text encoder (pretrained) for evaluation The No.
-of successes RN50 RN101 ViT-B/16 ViT-B/32 ViT-L/14 Ensemble BLIP-2 Clean image 0.609 0.662 0.646 0.651 0.506 0.615 381 MF-it (CLIP) [46] 0.606 0.658 0.643 0.648 0.503 0.612 370 MF-ii (CLIP) [46] 0.606 0.658 0.642 0.648 0.502 0.611 377 MF-it (EVA-CLIP) 0.650 0.698 0.684 0.689 0.554 0.655 547 MF-ii (EVA-CLIP) 0.646 0.696 0.679 0.685 0.545 0.650 529 InstructTA-MF-it (ours) 0.650 0.699 0.683 0.689 0.551 0.654 566 InstructTA-MF-ii (ours) 0.652 0.699 0.685 0.692 0.552 0.656 571
+Appendix 0.A Real Instructions
+Here are the 11 real instructions adapted from LLaVA-Instruct-150K:
+  • What’s your interpretation of what’s happening in this picture?
+  • What is taking place in this scenario?
+  • Describe the image’s visual elements with extensive detail.
+  • Could you explain the key elements of this picture to me?
+  • What are the essential components depicted in this image?
+  • What does this picture represent?
+  • Could you provide more detail on the components of the image shown?
+  • Compose an elaborate depiction of the presented picture.
+  • Examine the picture thoroughly and with attention to detail.
+  • Provide a description of the image below.
+  • What can you observe occurring in this picture?
+Table A6: Targeted attack performance with shuffled instructions.
+LVLM Attack method Text encoder (pretrained) for evaluation NoS ($\uparrow$) RN50 RN101 ViT-B/16 ViT-B/32 ViT-L/14 Ensemble BLIP-2 Clean image 0.275 0.397 0.319 0.328 0.177 0.299 2 MF-tt 0.277 0.400 0.320 0.331 0.177 0.301 3 MF-it 0.529 0.634 0.574 0.578 0.465 0.556 419 MF-ii 0.529 0.634 0.570 0.578 0.463 0.555 419 InstructTA 0.556 0.658 0.597 0.604 0.493 0.582 561
+Appendix 0.B Cross-Instruction Transferability
+To explore the cross-instruction transferability of targeted attacks, we shuffle the instruction-answer paired set given certain semantic similarities between the 11 instructions.
+The results on BLIP-2 are reported in Table A6.
+We can see that our InstructTA shows the best targeted attack performances in terms of CLIP-score and NoS.
+Hence, the adversarial examples generated by our method have superior cross-instruction transferability.
+Appendix 0.C Source Code
+The source code is available at https://github.com/xunguangwang/InstructTA.